OpenStudy (thatonegirl_):

calculus question! How do you know when to take the ln of an equation when finding a derivative?

OpenStudy (thatonegirl_):

For example... \[y=(sinx)^{x ^{2}}\] My teacher said to take the ln of both sides first but how would u know when to do tht?

zepdrix (zepdrix):

When you have a function raised to a function:\[\large\rm f(x)^{g(x)}\]When x is showing up in both the base and the exponent it creates a little bit of a problem. That's when logs really really help.

OpenStudy (thatonegirl_):

Oh okay thank you. So you couldn't just do chain rule for that?

satellite73 (satellite73):

"take the ln of both sides" lorda mercy

OpenStudy (thatonegirl_):

Lol i forgot the name xD

satellite73 (satellite73):

no no that is fine the truth is that the definition of \[b^x\]is \[e^{x\ln(b)}\] so the definition of \[(\sin(x))^{x^2}\] is \[\huge e^{x^2\ln(\sin(x)}\] and NOW you use the chain rule

zepdrix (zepdrix):

Consider constant value n for these next two examples: \(\large\rm f(x)^n\) <-- simply apply power rule into chain rule \(\large\rm \frac{d}{dx}f(x)^n=nf(x)^{n-1}\) \(\large\rm n^{f(x)}\) <-- apply exponential rule into chain rule \(\large\rm \frac{d}{dx}n^{f(x)}=n^{f(x)}(ln~n)f'(x)\) But when you x stuff in both the exponent and base, something else crazy is going on :)

zepdrix (zepdrix):

Sure sure sure, or you can do that little clever trick that Sat mentioned ^

satellite73 (satellite73):

@zepdrix not really a trick other way is trick to get the variable out of the exponent the way i wrote above appeals directly to the definition of an exponential

zepdrix (zepdrix):

Woops I didn't apply chain rule in my power rule :x soz

OpenStudy (thatonegirl_):

Ahh thank you both for the very thorough explanations :)

satellite73 (satellite73):

you will see that there is no free lunch, all the work entailf finding the derivative of \[x^2\ln(\sin(x))\] using the product and chain rule